find volume and surface area of prism to the nearest tenth.
H=2.4,B=3.7,L=3.7

Answer by ikleyn(52925)

.
find volume and surface area of prism to the nearest tenth.
H=2.4,B=3.7,L=3.7
--------------------------------------

Volume V = L*B*H = 3.7*3.7*2.4 = . . . (?)  
  Please calculate it yourself.


Surface area  S = 2*(LB + L*H + B*H) = 2*(3.7*3.7 + #.7*2.4 + 3.7*2.4) = . . . (?)
  Please calculate it yourself.